# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ 5 Robustness and Validation
We conduct several placebo tests to bolster the validity of our identification strategy and our theoretical framework. First, we argued that wives affected by their husbands’ WFH shocks are those with children of child-rearing age. Indeed, employment gaps are virtually non-existent between spouses that have no children, or whose children are old enough to be in school and/or take care of themselves. We can verify this hypothesis using the full CPS data. In Table B.22, we run a placebo test by estimating our baseline cross-section specification in the sample of households outside our restriction on the age of youngest child. Across the board, we verify that women’s monthly labor market outcomes are unaffected by their spouses’ WFH occupation-level shocks when their children are as old as teenagers (12 to 18 years old), are young adults (over 18 years old), or when they have no children at all.
Second, our preferred specification uses husbands’ changes in their occupation WFH rates following Covid-19 as an exogenous shock that impacts their wives’ labor market decisions. We expect that the nationwide WFH shocks only affect female spouses in households with young children, since they are usually the ones who bear the majority of child-rearing responsibilities. Men’s labor supply is close to one on average, as shown in the descriptive statistics from previous sections. In order to test this hypothesis, in Table B.23 we invert the estimation procedure and estimate the effect of female spouses’ WFH shocks on their husbands’ labor market outcomes using the monthly CPS data. By and large, we find no effect on men’s labor market outcomes, whether we look at women’s occupation and job industry shocks.
Third, we conduct a standard timing falsification test by coding the timing of the Covid shock as starting on March 2018 - three years before the pandemic began - and estimating the baseline specification on the 2017-2019 CPS subsample. This data has no influence from the post-Covid era, and since the Covid shock is largely unexpected, we would expect no effects to appear on these estimations. We present the results from this placebo test in Table B.24, which confirm that there are no significant employment, labor force participation or wage effects occurring around this artificial time of treatment.
Evidence from the ACS
We replicate our main results on wives’ labor market outcomes using the ACS sample (2017–2023), described in Section 2. While the ACS does not include all CPS outcomes, [^26] For example, it lacks a measure of annual labor force participation, and annual outcomes refer to the 12 months preceding the interview rather than the calendar year. We also observe only the interview year (not month), so we define the post-treatment period as 2022–2023 and include year fixed effects. it offers the advantage of a substantially larger sample size. Table B.25 confirms our main findings: treated wives experience a 2 percentage point increase in employment and labor force participation, work more hours per week, are 3 percentage points less likely to work part-time, and earn roughly 5% higher wages. When we examine heterogeneity by the age of the youngest child in Table B.26, we find that, much like the CPS results, the employment effects in the women ACS sample also appear to be driven by households with toddlers aged 0–4.
Evidence from the SIPP
Finally, we replicate several of our main employment results and analysis using the SIPP data. Panel A of Table B.27 shows that baseline estimates [^27] Since the SIPP only has data for reference years up until 2023, we moved the post-Covid period to after March 2022, and remove data between March 2020 and March 2022 for estimations. for the effect of husbands’ WFH occupation shocks on their wives’ labor market outcomes are positive in the SIPP cross-section. Moreover, point estimates using the SIPP data are larger than our CPS and ACS results. We are unable to test whether this difference is larger, since all estimates come from entirely different sources. Nevertheless, the size of the standard errors would suggest that the differences may not be statistically significant. In Panel B, point estimates suggest that employment effects are larger for households with younger children - in the range of 0 to 4 years old - but the differences are not statistically significant.
As with the CPS and ACS, we perform a heterogeneity analysis based on the age of the youngest child, separately for households with multiple children and with only one children. These estimations are shown in Table B.29, and overall these are closely aligned with the CPS heterogeneity results: We find some evidence that households with toddlers experienced significant effects, whereas non-toddler households experienced non-significant effects - as seen by the test on the sum of the coefficients. This is particularly true in households with more than one child. However, because the interaction coefficients are not significant, we cannot interpret this as conclusive evidence of heterogeneous results.
In addition to our cross-section specification, the panel component of the SIPP data allows us to estimate a panel regression. Specifically, we replace our baseline fixed effects with month-year and household-level fixed effects. We present results for several samples in Table B.28.282828In addition to sampling households of dual parent households with children aged 0 to 8 years old, we restrict the sample to households that have not switched treatment status across different panel interviews. In Panel A, which uses all data from the 2017 to 2023 reference years, we find positive and large estimates for employment, labor force participation, and wage effects for women; however, these estimates are not statistically significant except for wages. The results similarly hold when we use data from the 2020 SIPP panel, which follows households from reference years 2019 to 2022, and thus is the only panel to cover same household before March 2020 and after March 2021.292929Due to logistical issues, the SIPP did not follow through the sample of new households that joined the sample in SIPP 2019. In panels C and D, we estimate our sample on balanced 2019-2022 and 2019-2021 panels, respectively. [^30] For panel D we change the post-Covid window to data after March 2021. Estimates are non-significant, which we argue is due to loss in power as we only observe a very limited number of different households in the balanced panels.
In Table B.30, we replicate and consolidate our main placebo and falsification tests developed for the CPS sample in our SIPP sample. These effects include estimating the effects of women’s WFH occupation shocks on their male spouses’ labor market outcomes, estimating women’s labor market effects on a sample without children or whose youngest child are teenagers, and moving the Covid-19 shock to March 2018 for a falsification test. The falsification test for households without children and the artificial timing of Covid-19 (Panels C and D) show the results we expected. The latter is specially reassuring as it provides evidence against pre-trends in the SIPP data. The labor employment effects for husbands (Panel A) are weakly significant and less than a half of the treatment effect for wives in the SIPP sample. However, women whose youngest are teenagers-to-adults see large effects, which goes against our expectations.
@@
Taken together, these robustness and falsification checks support the validity of our identification strategy. The estimated effects are not driven by differential selection, sample restrictions, or alternative treatment definitions. Main estimates across different nationally representative datasets are similar in magnitude and sign.
6 Mechanisms
Understanding the household mechanisms that enabled women to increase their labor supply when their husbands became more likely to work from home is central to interpreting the overall effects of the WFH shock. We posit that our large employment effects for women with young children, to some extent, can be explained if we also observe changes in the household production function, and more specifically, in the burden of child-rearing. This could occur through the following channels: (i) a substitution channel, where greater time spent at home of the husband leads to an increase of his involvement in childcare, potentially substituting not only for the wife’s time but also for alternatives such as paid daycare; or (ii) a multitasking channel, where women provide childcare while simultaneously engaging in other activities, such as working from home or performing household tasks, effectively changing their childcare “production function”.
To explore how the intra-household allocation of time adjusts in response to the WFH shock, we estimate Equation (1) using variables from the ATUS and SIPP that capture both the intensive and extensive margins of labor supply and household production. [^31] Since the ATUS and SIPP have fewer observations, we moved the post-Covid period to after March 2022, and remove data between March 2020 and March 2022 for estimations. We still lose a lot of precision due to sample size. Our goal is to assess how time spent on domestic tasks and childcare is redistributed within households, enabling the observed increase in labor market participation among women in treated households. [^32] For the ATUS, we restrict the sample to respondents who completed the time-use diary on a weekday and non-holiday day, in order to capture household dynamics on a typical workday.
Time-Use in ATUS
First, the increase in time allocated to market work is consistent across spouses and aligns with the labor supply effects documented in Section 4. As shown in Table 3, treated women increase the share of their day spent working by 7.5 percentage points, i.e., equivalent to nearly two additional hours. [^33] A full day includes 24 hours This rise is driven entirely by an increase in main job hours, with significant gains only from working from home. Although the estimate for time spent working outside the home is not statistically significant, both its direction and the pattern shown in Figure A.7 suggest an increase during the first two years of the post-treatment period. We find no significant changes in time spent commuting or in hours worked at a secondary job.
@@ 6 Mechanisms
In contrast to their wives, husbands experience a substantial reallocation in their work location, consistent with being the directly treated group. They increase time working from home by 4.7 percentage points and appear to reduce time working outside the home. Although the latter estimate is noisy due to limited sample size, Figure A.7 shows a negative and relatively stable effect across all post-treatment years. These shifts in work location are persistent over time, unlike the more transitory patterns observed for women. Notably, men do not exhibit any change in total hours worked. However, treated men consistently report reduced commuting time, a pattern strongly aligned with the nature of the WFH shock, and an increase in time spent on secondary jobs.
Second, treated households exhibit meaningful shifts in the allocation of household and childcare responsibilities. On the household production side, women spend significantly less time shopping, while time spent on general household tasks such as cooking and cleaning shows no statistically significant change. Although these estimates are not precise, they consistently point in a negative direction, as shown in Figure A.8. Regarding childcare, although total time spent caring for children does not change significantly, there is a clear shift in how childcare is provided. Women in treated households reduce the amount of time spent on childcare as a primary activity by 2.7 pp, while not changing the time spent on childcare while engaged in other primary non-work activities. This suggests that women are more likely to work and to work longer hours, and one way to cope with this is to reduce the time spent on childcare as the main activity. As a result, secondary childcare becomes more prominent, with its share of total childcare time increasing by 8.4 pp relative to a pre-treatment mean of 64.5 percent.
Husbands, on average, do not exhibit significant changes in total time spent on household production or overall childcare. However, as shown in Figure A.8, their time spent on primary childcare increases significantly in most post-treatment years, while time spent on secondary childcare remains unchanged. In particular, treated husbands are no more likely to report engaging in childcare while working, which helps explain the decline in secondary childcare as a share of their overall caregiving. These patterns point to a key mechanism: as women reduce time spent on primary childcare, men appear to partially compensate—a finding consistent with the substitution channel. That said, the extent of this compensatory behavior is limited and does not fully offset the reductions observed among women.

### v2

@@ 6 Robustness and Validation
We conduct a standard timing falsification test by artificially assigning the onset of the COVID shock to 2018 and estimating the baseline specification on the 2017–2020 ACS subsample, with reference years 2016–2019. We define the pre- and post-shock periods as 2016–2017 and 2018–2019, respectively. Because this subsample predates the COVID period and the shock was largely unanticipated, we would not expect to observe treatment effects in this placebo setting. Results from this falsification exercise are reported in Table B.31. We find no evidence of effects on earnings, hours worked, weeks worked, or part-time status around this artificial treatment date. We do observe a small effect on log wages, as well as on employment, which are consistent with the mild pre-trends and patterns documented in the dynamic DiD plots.
Evidence from Alternative Datasets
Estimate from alternative datasets generally support our results. First, we replicate our analysis using the 2017-2024 CPS sample, as described in Section 3. Although the CPS offers a smaller sample than the ACS, it provides comparable labor-market measures that allow an independent check of our main findings. [^41] Because CPS outcomes are measured monthly, we define the “during” and “post” periods as described in Section 3 for the ATUS sample. Namely, we define March 2020 as the starting point of the post-treatment period and March 2022 as the end date of the during period. Consistent with this, Table B.32 reports reduced-form estimates for wives that closely mirror the ACS results. Women in higher-exposure households are 2.7–3.5 percentage points more likely to be employed, depending on whether employment is measured over the previous week or year. They report higher weekly hours, are roughly 3 percentage points less likely to work part-time, and earn approximately 8 percent higher wages.
Second, we replicate several of our main employment results for women using SIPP data and exploit its four-year panel structure to estimate panel regressions. Specifically, we replace the baseline fixed effects with month–year and household fixed effects. Results for several samples are reported in Table B.33.424242In addition to restricting the sample to dual-parent households with children aged 0–8, we further limit the analysis to households that do not switch treatment status across panel interviews. Although the SIPP contains information from 2017 to 2023, we focus on the 2020 SIPP panel, which follows households from reference years 2019 to 2022 and is the only panel spanning both the pre- and post-COVID periods. [^43] Due to logistical constraints, the SIPP did not continue following households newly added in the 2019 sample.
Across specifications, the estimated effects of spousal WFH shocks on women’s labor supply have the same sign as in the ACS and CPS, though the estimates are substantially noisier. Extensive-margin effects are similar in magnitude but less precisely estimated, while hours worked and earnings effects are statistically significant but larger than those observed in the other datasets. We interpret these inflated magnitudes cautiously, as they are likely driven by limited statistical power: the panel analysis is based on a much smaller sample, with approximately 500 unique households.
@@
Taken together, these robustness and falsification checks support the validity of our identification strategy. The estimated effects are not driven by differential selection, sample restrictions, or alternative treatment definitions. Main estimates across different nationally representative datasets are similar either sign, although panel estimates are noisy.
7 Household Specialization
The remote work shock may have changed how households coordinate their market and home production decisions. Understanding and evaluating the changes in household specialization and time use is central to reconcile not only women’s increase labor supply, but also interpret the overall effect of the WFH shock on households.
In this section, we argue that the large employment effects we observe for women with young children can be explained by changes in the household production function, particularly in the burden and organization of childcare duties. As discussed in Section 5, these changes may arise from two forces: (i) spillovers from husbands’ increased presence at home and (ii) women’s own access to work from home through assortative matching.
To explore how the intra-household allocation of time adjusts in response to the WFH shock, we estimate Equation (4) using variables from the ATUS, ACS and SIPP that capture both the intensive and extensive margins of labor supply and household production. Our goal is to assess how time spent on household activities and childcare is redistributed within households, explaining with the lens of a collective perspective the observed increase in labor market participation among women in treated households. [^44] For the ATUS, we restrict the sample to respondents who completed the time-use diary on a weekday and non-holiday day, in order to capture household dynamics on a typical workday. [^45] We evaluate the comparability between the two ATUS samples and the CPS, given that the ATUS samples are non-overlapping subsamples drawn from the CPS. Table B.38 presents the first and second moments of the distribution for key characteristics of husbands, wives, and households across the CPS sample and both ATUS subsamples. The CPS sample displays modest differences relative to the ATUS samples, particularly in the likelihood of treatment and in measures of household income and education. These patterns suggest that the ATUS subsamples tend to capture relatively wealthier and more educated households. [^46] The estimated effects might be driven by specific subsamples whose responses to the WFH shock diverge from those identified in the ACS results in Table 2 or the CPS results in Table B.32. We address this in two ways. First, both Table B.34 and Figure 4 show that the increase in women’s employment following the WFH shock is statistically significant and persistent. Second, despite the smaller sample sizes in ATUS (mainly on the ASEC variables like wages, earnings and weeks worked), Table B.39 shows that the estimated effects on wives’ labor market outcomes are directionally consistent with those in CPS sample, which are also aligned with our main findings using the ACS sample.
Market Production Time Use
@@ 7 Household Specialization
First, husbands experience a substantial change in their work location, consistent with being the directly treated group as Figure A.10 illustrates. These sheds more light on the validity of the first stage. They increase time working from home by 4.7 percentage points and appear to reduce time working outside the home. Although the latter estimate is noisy due to limited sample size, Figure 4 shows a negative and relatively stable effect across all post-treatment years. These shifts in work location are persistent over time, unlike the more transitory patterns observed for women. Notably, men do not exhibit any change in total hours worked. However, treated men consistently report reduced commuting time by 5 minutes on average. Efficiency-wise these results mean that men are working the same amount of time, but avoiding part of the sunk costs from commuting. [^47] For the ATUS sample, the reduction is of around 17 minutes.
Wives experience an increase in the time allocated to market work, which aligns with the labor supply effects Section 5 documents. As shown in Figure A.10, treated women increase on average the share of their day spent working after the pandemic by 7.5 percentage points, i.e., equivalent to nearly two additional hours. [^48] A full day includes 24 hours. This rise is driven mainly by a 5 percentage point increase in working from home, which is consistent with the 6.5 pp. increase in WFH uptake we report in Figure 4. Consistent with the fact that most of increase in working is coming from work from home uptake, there is a significant reduction in the time spent commuting for most of the post-treatment years by almost 10% - 2.5 minutes. [^49] In Table B.34, the ATUS commuting variable is not significant, which it is not striking. It is more difficult to detect an effect on a substantially smaller sample size and on an outcome whose variation is small, relative to others like share of time working. By performing the disentanglement of spillovers from correlated direct effects as in Section 5, we observe that most of women’s increase in WFH uptake and reduction in commuting comes from the correlated direct effects, as shown in Table B.35. This result is alligned with the assumption proposed in the extension of Section 2 in which women can also hold remote work jobs.
Household Production Time Use
This change in the joint labor market choices came alongside with new patterns in the collective household production, as Figure 4 documents. Treated households exhibit meaningful shifts in the allocation of childcare responsibilities.
Although the total time spent caring for children does not change significantly, there is a clear shift in how childcare is provided. Women in treated households reduce the amount of time spent on childcare as a primary activity by 2.7 pp - an approximately 25 percent decrease relative to the average pre-treatment values. While they are not reporting to be spending significant more time child-caring while working, although they are spending more time WFH. Additionally, there is no reduction in secondary childcare, as Table B.34 shows. In other words, women mainly cut back on primary childcare. As a result, secondary childcare becomes more prominent: its share of total childcare time rises by 8.4 p.p. relative to a pre-treatment mean of 64.5 percent.
Husbands, on average, do not exhibit significant changes in total time spent on overall childcare. However, as shown in Figure 4, their time spent on childcaring while working increases significantly by 1.5 percentage points. This is a 200% increase relative to baseline. This aligns with the fact that men are more likely to work from home due to their shock. These patterns point that women reduce time spent on primary childcare, and men appear to partially compensate while working. In terms of household division, there is a more equal distribution in the time spent childcaring between spouses. The gender gap in childcare time was of approximately 0.5 in the pre-treatment, as Table B.34 reflects, meaning that women spent the double of time childcaring compared to men. Treated households narrow the gap to approximately 0.75.505050On the household production side, both men and women do not spend significantly any change in time in housework. Although this estimate is not precise for women, it consistently points in a negative direction, while virtually zero for men. Childcare might be more substitutable and proportionally more important in time to be able to detect significant changes, compared to other household tasks. This might explain the lack of precise effects on the other less prominent household activities.  [^51] We conduct a standard placebo test by assigning the onset of the post-treatment period to March 2018. Table B.40 reports the results. We find no significant changes.
Figure 4: Dynamic Effect of Husbands’ WFH Shock on Women’s WFH Uptake and Households’ Time-Use
(a) Women WFH Uptake (ACS)
Women Men
(b) Commuting (ACS)
Refer to caption Refer to caption
(c) Primary Childcare (ATUS)
Refer to caption Refer to caption
(d) Childcare while Working (ATUS)
Refer to caption Refer to caption
Notes: The top center panel shows effect of spousal WFH shock on women’s WFH uptake. Starting from the second row of subfigures, each panel to the left plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. Panels to the right compare husbands directly. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and based on standard errors clustered at the husband’s occupation level.
Childcare in SIPP
Using the SIPP childcare topical modules, we are able to provide complementary evidence how household change their sources of childcare and hiring of daycare following shocks on spousal WFH options. We present some results in Table B.36. We estimate our baseline cross-section specification on the 2018-2024 SIPP panel datasets on a battery of childcare related variables. Our results show that in households shocked by the occupation-level WFH treatment, when the child’s mother is working, the spouse is 6.4 pp more likely to be involved in childcare, or roughly a 13% increase over the baseline. On the other hand, the woman herself is not any more likely to be involved in childcare.
We also find that households exposed to the WFH shock are significantly less likely to rely on grandparents or other relatives, or on daycare, for day-to-day childcare. Treated households are 7.3 pp less likely to have any paid childcare, and although imprecisely estimated, their total childcare expenditure falls by about 16%. The dynamic DiD estimates in Figure A.11 show clear post-2019 declines in both the probability of paying for childcare and log childcare spending. Overall, these patterns are consistent with fathers becoming more available for childcare in treated households.
Household’s Well-being